OpenStudy (anonymous):

2.  What is a solution of x^2 + 6x = –5? x = –6 x = –1 x = 1 x = 6

OpenStudy (inkyvoyd):

x^2 + 6x +5=0 Factor.

OpenStudy (anonymous):

I'll try.

OpenStudy (anonymous):

x^2 + 6x = –5 x^2 + 6x +5= 0 (x+1)*(x+5)=0 x=-1 x=-5

OpenStudy (anonymous):

I got -1.

OpenStudy (inkyvoyd):

@anupam1031 , nice sniping.

OpenStudy (anonymous):

lol

OpenStudy (inkyvoyd):

Feo, did you factor?

OpenStudy (anonymous):

Its B lol

OpenStudy (anonymous):

thanks yall, appreciate it.